### v3

# CLIP-EBC: CLIP Can Count Accurately through Enhanced Blockwise Classification

Abstract
We propose CLIP-EBC, the first fully CLIP-based model designed for accurate crowd density estimation. Existing classification-based crowd counting frameworks face challenges when directly applied with CLIP. For instance, these methods quantize count values into bordering real-valued bins, which are inconsistent with CLIP’s pretraining corpus. Besides, this quantization strategy also introduces label ambiguity near shared bin borders. To address these issues, we propose the Enhanced Blockwise Classification (EBC) framework, which utilizes integer-valued bins and additionally incorporates a density-map-based loss to further improve count accuracy. Building on this backbone-agnostic framework, CLIP-EBC fully leverages CLIP’s recognition capabilities for crowd counting. Experiments show that EBC improves classification-based methods by up to 44.5% in mean absolute error (MAE) on the UCF-QNRF dataset. Furthermore, CLIP-EBC achieves state-of-the-art performance on the NWPU-Crowd test set, with an MAE of 58.2, surpassing the previous best method by 8.6%. Our code is available at https://github.com/Yiming-M/CLIP-EBC.
Index Terms:
crowd counting, crowd density estimation, CLIP
I Introduction
The vision-language model CLIP [1] is pre-trained on millions of image-text pairs using contrastive learning, enabling it to align visual and textual modalities effectively. This capability has allowed CLIP to excel in pattern recognition, setting new benchmarks in downstream tasks such as zero-shot image classification [1], and semantic segmentation [2]. Despite its strong performance in these recognition-based tasks, its potential for regression tasks, which require predicting continuous values, remains largely unexplored. This limitation arises primarily from its contrastive learning objective, which is inherently designed for discrete decision-making tasks, such as classification or retrieval, where the goal is to match or rank embeddings. However, regression tasks require predicting continuous values, which is not the focus of contrastive learning.
In this work, we aim to address this gap by investigating whether CLIP can effectively perform regression tasks, with a specific focus on crowd counting. Crowd counting is a particularly demanding task due to the inherent challenges, such as high-density crowds, occlusions, and scale variations. Thus, it provides a robust testbed to evaluate CLIP’s applicability to regression problems.
Figure 1: Comparison of our model CLIP-EBC and other CLIP-based crowd counting methods. Top: CrowdCLIP [3] predicts total counts from global image features. Due to the incapability to capture spatial details, it suffers from reduced accuracy. Middle: CLIP-Count [4] employs extra modules to generate density maps, improving accuracy at the cost of increased complexity. Bottom: CLIP-EBC is the first fully CLIP-based density-map approach, achieving high accuracy without additional modules.
Crowd counting is typically formulated as a blockwise density map regression task [5, 6, 7]. Therefore, an intuitive approach to adapt CLIP for crowd counting is through classification frameworks, where blockwise count values are quantized into discrete bins. However, directly applying CLIP in existing classification frameworks [8, 9] presents significant challenges.
Figure 2: Overview of the EBC framework and the CLIP-EBC model. The EBC framework introduces an integer-based quantization strategy and noise reduction to partition count values into $n$ predefined bins (e.g., $\{0\},\{1\},\cdots,[m,\infty)$). Given an input image, the crowd counting model implemented under EBC generates a blockwise probability map. The density map is obtained by averaging the bin midpoints weighted by the probability map. Finally, the predicted probability map and the density map are fed into the proposed $\mathcal{L}_{\text{DACE}}$ loss (Eq. (III-A4)), which is used to fine-tune the model. CLIP-EBC, an instance of EBC, leverages the predefined bins and the prompt template to construct $n$ text prompts (e.g., “There is 0 person”). The frozen CLIP [1] text encoder extracts text embeddings, while the fine-tuned CLIP image encoder generates image feature maps. For the image feature vector at each spatial location, we calculate its cosine similarity with the text embedding for each bin and then apply softmax to obtain the probability map. Subsequent steps align with the EBC framework to generate the final outputs.
For instance, existing classification methods employ Gaussian smoothing to reduce the sparsity of the ground-truth density maps, transforming the initially discrete count value spaces $\mathbb{N}$ into a continuous space $\mathbb{R}_{+}$. This requires dividing the space into consecutive real-valued intervals to construct classification bins. While such bins can cover $\mathbb{R}_{+}$, they are poorly suited for constructing CLIP-compatible text prompts, since CLIP’s text encoder is pre-trained on natural language corpora, where descriptions are typically human-readable and discrete (e.g., “there are 3 people”). Prompts involving real values (e.g., “there are 3.14 people”) lack semantic clarity and may result in noisy text embeddings and degraded performance. Besides, the use of adjacent bins introduces label ambiguity, particularly for samples near shared boundaries. For example, the bins [0, 0.05) and [0.05, 0.1) in [8] create uncertainty for a block with a count of 0.05, as it is unclear which bin it should belong to. Such ambiguity weakens the reliability of predictions. Another limitation of current classification-based methods is their sole focus on optimizing the classification accuracy. This can lead to suboptimal performance on regression metrics, such as Mean Absolute Error (MAE) or Root Mean Squared Error (RMSE), which are critical for counting tasks.
To address these challenges, we propose the Enhanced Blockwise Classification (EBC) framework, specifically designed to overcome the limitations of current classification-based methods. Our EBC framework is model-agnostic: existing regression-based methods can be implemented using EBC with minimal modifications to achieve substantial performance improvements. Building on this framework, we introduce CLIP-EBC, which fully leverages the original structure of CLIP for crowd counting. Unlike other CLIP-based methods [3, 4] (see Fig. 1), CLIP-EBC is the first model to fully utilize CLIP for density-map-based crowd counting. While CrowdCLIP [3] also relies solely on CLIP, it predicts total counts instead of density maps, limiting its ability to capture fine-grained spatial details for accurate predictions.
Experiments across four benchmark datasets demonstrate the significant advantages of our EBC framework. For instance, EBC reduces Mean Absolute Error (MAE) by up to 44.5% on the UCF-QNRF dataset [10] compared to the standard blockwise classification method [8]. Furthermore, CLIP-EBC surpasses state-of-the-art crowd counting methods, achieving an MAE of 58.2 and a Root Mean Square Error (RMSE) of 268.5 on the NWPU-Crowd test split [11]. These results demonstrate that CLIP, with the support of the EBC framework, can accurately estimate crowd density distributions. To summarize, our contributions are as follows:
  • We propose the novel Enhanced Blockwise Classification (EBC) framework that substantially improves over the classification-based methods by addressing issues related to Gaussian smoothing and loss functions.
  • Building upon EBC, we present the first fully CLIP-based crowd counting model CLIP-EBC that is capable of accurately estimating crowd densities.
  • We conduct extensive experiments across multiple data-sets to showcase the effectiveness of EBC in enhancing existing methods and the competence of CLIP-EBC as a state-of-the-art crowd-counting method.
II Related Work
Classification-Based Crowd Counting Methods. These models aim to address the long-tail distribution of count values, where large counts are underrepresented, hindering the performance of regression-based models. These methods partition the range $[0,\infty)$ into consecutive non-overlapping intervals to increase the sample size per class. For instance, Xiong et al.[9] proposed DCNet, which predicts counts at multiple levels using a shared set of bins. However, this approach fails to account for the lower likelihood of large counts at lower levels, exacerbating class imbalance. Liu et al.[8] addressed this by introducing blockwise classification, where models output probability maps. However, these methods rely on Gaussian smoothing, which introduces label noise and boundary ambiguities, leading to reduced accuracy. Also, optimizing for classification accuracy alone cannot guarantee optimal performance under regression metrics such as MAE.
CLIP in Crowd Counting. CLIP [1] has demonstrated exceptional performance in recognition-related tasks, such as zero-shot image classification [1], and semantic segmentation [2], but there are only few studies in regression-related tasks, particularly crowd counting. Liang et al.[3] proposed CrowdCLIP, an unsupervised (or weakly-supervised) fully CLIP-based model, which operates on global image features extracted by CLIP’s image encoder. This global representation lacks spatial granularity and cannot capture the fine details or density variations across different parts of the image. Jiang et al.[4] extended CLIP for text-guided zero-shot object counting, incorporating additional modules such as multimodal interaction layers and decoders to generate density maps. In contrast, we propose CLIP-EBC, the first crowd counting framework fully based on CLIP, designed to estimate crowd density maps without relying on external modules. Fig. 1 presents the comparison with these methods .
III Methods
Our Enhanced Blockwise Classification (EBC) framework improves the Standard Blockwise Classification (SBC) [8] from three key aspects: the quantization strategy, noise control, and loss formulation. In this section, we first briefly recapitulate SBC, and then provide a detailed explanation of our enhancements. Finally, we demonstrate how to leverage our model-agnostic EBC framework with CLIP for density map estimation, introducing our model CLIP-EBC. Fig. 2 illustrates the overview of EBC and CLIP-EBC.
III-A Enhanced Blockwise Classification
III-A1 Recap: Standard Blockwise Classification (SBC)
Similar to many regression-based methods [5, 8, 6, 7], SBC also employs blockwise prediction. Formally, given an input image $\boldsymbol{X}\in\mathbb{R}_{+}^{3\times H\times W}$, with $3$ channels, height $H$ and width $W$, SBC outputs a probability map
$\boldsymbol{P}^{*}\in\mathbb{R}_{+}^{(n,\,H//r,\,W//r)},$ (1)
where ///// / represents the floor division operator and the integer $r$ is a model-related reduction factor. At each spatial location $(i,\,j)$, the $n$-dimensional vector $\boldsymbol{P}^{*}_{:,\,i,\,j}$ represents the predicted probability distribution of the count value in the block $(r(i-1):ri,\,r(j-1):rj)$ of the image over the $n$ predefined bins $\{\mathcal{B}_{i}\mid i=1,\cdots,n\}$. These bins satisfy
$\forall p\neq q,\,\mathcal{B}_{p}\cap\mathcal{B}_{q}=\emptyset\quad\text{and}\quad\mathcal{S}\subseteq\cup_{i=1}^{n}\mathcal{B}_{i},$
where $\mathcal{S}$ denotes the set of all possible count values. The predicted density map $\boldsymbol{Y}^{*}$ can be generated by taking the expectation or the mode over the bins. In the former case, $\boldsymbol{Y}^{*}$ is calculated by
$\boldsymbol{Y}^{*}_{i,j}=\sum_{k=1}^{n}\boldsymbol{P}^{*}_{k,i,j}\cdot b_{k},$ (2)
where $b_{k}$ is the average count value in the bin $\mathcal{B}_{k}$.
[…]
III-A2 Enhancement 1: Integer-Valued Bins
[…]
(a) Region with noisy labels
(b) Zoomed-in view of the region
Figure 3: An example from ShanghaiTech A dataset [12] illustrating an extremely dense area with erroneous annotations. The $8\times 8$ magenta box in Fig. 3(a) highlights a congested region with an labeled count of 9 individuals. However, the zoomed-in view in Fig. 3(b) reveals that no discernible human figures can be identified within the marked box, evidencing label noise in highly dense regions.
Integer-Valued Bins
Noise Reduction
Combined Loss
MAE $\downarrow$
140.6
88.3
85.8
✓($\lambda=0.01$) 83.9
✓($\lambda=0.10$) 83.4
✓($\lambda=1.00$) 77.9
✓($\lambda=2.00$) 82.3
TABLE I: Influence of each enhancement in our EBC framework evaluated on UCF-QNRF [10]. When no enhancement is utilized, the model deteriorates into SBC [8]. The reduced mean absolute errors demonstrate the effectiveness of our enhancement strategies and the superiority of EBC.
III-A3 Enhancement 2: Noise Reduction
Another challenge often overlooked by existing approaches is the presence of erroneous and noisy annotations within highly dense crowd areas, as illustrated in Fig. 3. This issue can stem from two primary factors: 1) low-resolution images, where annotators struggle to precisely determine the head counts in congested areas; and 2) datasets with images resized to significantly smaller resolutions after labeling to optimize storage and training time. For instance, consider an HD image annotated with 1000 people, but subsequently resized to $64\times 64$ pixels. In such a scenario, each individual occupies an average of only 4 pixels––insufficient to distinctly represent a human figure. To address this problem, we propose a noise reduction strategy by merging all count values equal to or greater than a preset threshold $m$ into a single large bin. This approach is integrated into our three binning strategies as follows:
Fine: $\,\{0\},\{1\},\cdots,[m,\infty)$ (3)
Coarse: $\,\{0\},\{1,2\},\cdots,[m,\infty)$ (4)
Dynamic: $\,\{0\},\cdots,\{n\},\{n+1,n+2\},\cdots,[m,\infty)$ (5)
An additional advantage of this strategy is that it increases the sample size of large count values by merging them into one bin, thereby alleviating the long-tail distribution issue.
III-A4 Enhancement 3: Combined Loss.
Previous blockwise classification methods [8, 9] primarily focus on minimizing the cross-entropy loss between the predicted probability map $\boldsymbol{P}^{*}$ and the ground truth $\boldsymbol{P}$ but often overlook the mismatch in count values. Since two different probability distributions can yield the same classification error while having different expectations (or modes) for the predicted count values, models trained with such loss functions may not guarantee accurate count estimates. To address this limitation, we propose the Distance-Aware-Cross-Entropy (DACE) loss:
$\mathcal{L}_{\text{DACE}}=$ $\mathcal{L}_{\text{class}}(\boldsymbol{P}^{*},\,\boldsymbol{P})+\lambda\mathcal{L}_{\text{count}}(\boldsymbol{Y}^{*},\,\boldsymbol{Y})$
$=$ $-\sum_{i=1}^{H//r}\sum_{j=1}^{W//r}\sum_{k=1}^{n}\boldsymbol{1}(\boldsymbol{P}_{k,i,j}=1)\log\boldsymbol{P}^{*}_{k,i,j}$
$+\lambda\mathcal{L}_{\text{count}}(\boldsymbol{Y}^{*},\,\boldsymbol{Y}),$ (6)
where $\boldsymbol{1}$ is the indicator function, $\boldsymbol{P}$ is the one-hot encoded ground-truth probability map, $\boldsymbol{P}^{*}$ is the predicted probability map, $\boldsymbol{Y}$ is the ground-truth density map, $\boldsymbol{Y}^{*}$ is the predicted density map, and $\lambda$ is a weighting factor balancing $\mathcal{L}_{\text{class}}$ and $\mathcal{L}_{\text{count}}$. Since the ground-truth density maps are not smoothed, we use the DMCount [6] based on the discrete Wasserstein distance as the count loss $\mathcal{L}_{\text{count}}$.
Backbone Type
MAE
MAE
Improvement
VGG16 C 140.6 77.9 44.5%
ResNet101 C 127.3 79.7 37.3%
MobileNetV2 C 135.9 83.2 38.7%
DenseNet201 C 118.1 82.9 29.8%
CSRNet R 119.2 79.3 33.4%
DMCount R 85.6 77.2 9.8%
TABLE II: Our EBC framework can improve both classifica-tion-based and regression-based methods. We reimplement both SBC models [8] and two regression-based methods (CSRNet [5] and DMCount [6]) using our EBC framework on the UCF-QNRF dataset. Regardless of the choice of the backbone, our EBC always improves the performance.
III-B The Structure of CLIP-EBC
As CLIP-EBC is implemented under our proposed EBC framework, here we focus on how to generate the predicted probability map $\boldsymbol{P}^{*}$. Additional essential details, such as the loss function used in training, are provided in Section III-A.
The image encoder of CLIP-EBC comprises a feature extractor $\mathcal{F}$, an interpolation layer $\mathcal{U}$, and a projection layer $\mathcal{P}$. We utilize CLIP’s image encoder without the final pooling and projection operations to extract the feature map
$\boldsymbol{H}=\mathcal{F}(\boldsymbol{X})\in\mathbb{R}^{c\times(H//p)\times{W//p}},$ (7)
where $c$ represents the number of output channels (for the ResNet backbone) or the embedding dimension (for the ViT backbone), and $p$ represents the model reduction factor (ResNet) or the patch size (ViT). Subsequently, $\boldsymbol{H}$ is passed to $\mathcal{U}$ to adjust the spatial size
$\tilde{\boldsymbol{H}}=\mathcal{U}(\boldsymbol{H})\in\mathbb{R}^{c\times(H//r)\times{W//r}}.$ (8)
Finally, we employ one $1\times 1$ convolutional layer $\mathcal{P}$ to project $\tilde{\boldsymbol{H}}$ into the CLIP embedding space, yielding
$\boldsymbol{F}^{\text{img}}=\mathcal{P}(\tilde{\boldsymbol{H}})\in\mathbb{R}^{d\times(H//r)\times{W//r}},$ (9)
where $d$ is the dimension of the CLIP embedding space.
Methods Type ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU (Val)
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
CrowdCLIP [3] U 146.1 236.3 69.3 85.8 283.3 488.7 - -
CLIP-Count [4] R 192.6 308.4 45.7 77.4 - - - -
DMCount [6] R 59.7 95.7 7.4 11.8 85.6 148.3 70.5 357.6
P2PNet [13] R 52.7 85.0 6.3 9.9 85.3 154.5 77.4 362.0
CLTR [14] R 56.9 95.2 6.5 10.2 85.8 141.3 61.9 246.3
STEERER [15] R 54.5 86.9 5.8 8.5 74.3 128.3 - -
S-DCNet [9] C 58.3 95.0 6.7 10.7 104.4 176.1 - -
UEPNet [16] C 54.6 91.2 6.4 10.9 81.1 131.6 - -
CLIP-EBC (ResNet50, ours) C 54.0 83.2 6.0 10.1 80.5 136.6 38.6 90.3
CLIP-EBC (ViT/B-16, ours) C 52.5 85.9 6.6 10.5 80.3 139.3 36.6 81.7
TABLE III: Comparison of our model CLIP-EBC with recent crowd counting approaches on ShanghaiTech A & B [12], UCF-QNRF [10] and NWPU-Crowd (Val) [11]. The “Type” column groups all methods into unsupervised (“U”), regression-based (“R”), or classification-based (“C”). The best results are highlighted in bold, and the second best results are underlined. Our model CLIP-EBC demonstrates competitive performance across all datasets, with particularly significant superiority on the NWPU validation split, where it outperforms all existing methods. : Results of cross-dataset evaluation.
For text feature extraction, we start by generating the input text prompts. Given a set of bins $\{\mathcal{B}_{i}\mid i=1,\cdots,n\}$, we create a text prompt for each bin $\mathcal{B}_{i}$ according to the following rules:
  • If $\mathcal{B}_{i}=\{b_{i}\}$, the text prompt is “There is/are $b_{i}$ person/people”, with the choice between “is/are” and “person/people” determined by whether $b_{i}>1$.
  • If $\mathcal{B}_{i}$ is finite and contains more than one element, then the text prompt becomes “There is/are between $\min(\mathcal{B}_{i})$ and $\max(\mathcal{B}_{i})$ person/people”. The decision between “is/are” and “person/people” is again made to ensure grammatical correctness.
  • If $\mathcal{B}_{i}=[m,\infty)$, then the text prompt is “There are more than $m$ people”.
Next, the resulting $n$ text prompts are tokenized using CLIP’s tokenizer. We then input the tokenized text into the original CLIP text encoder, with its parameters frozen during training. This process yields text features $\boldsymbol{F}^{\text{txt}}$ with dimensions $d\times n$. The probability map $\boldsymbol{P}^{*}$ can be computed from the image feature map $\boldsymbol{F}^{\text{img}}$ and the text features $\boldsymbol{F}^{\text{txt}}$. First, we calculate the cosine similarity between the image feature vector $\boldsymbol{F}^{\text{img}}_{:,\,i,\,j}$ and the $n$ extracted text embeddings. We then normalize these similarities using softmax to obtain $\boldsymbol{P}^{*}_{:,\,i,\,j}$, which denotes the probability scores across the $n$ bins for block $(i,\,j)$. To obtain the predicted density map $\boldsymbol{Y}^{*}$, we use Eq. (2). We train CLIP-EBC with our loss function in Eq. (III-A4).
IV Experiments
Implementation Details. We set the block size $r=8$ (unless explicitly specified), which is commonly used in many works [5, 8, 6]. For this block size, we set $m=4$. We use the Adam optimizer [17] to train all our models with an initial learning rate of $41e-$ 4, which is adjusted through a cosine annealing schedule. The batch size is fixed at $8$ for all datasets. For our ViT-based CLIP-EBC model, we apply visual prompt tuning [18] by prepending 32 learnable tokens to each layer, rather than fine-tuning the entire backbone. Following existing methods, we use the mean absolute error (MAE) and root mean square error (RMSE) to evaluate our models.
IV-A Effectiveness of EBC
We first demonstrate the effectiveness of our EBC framework using VGG16 [19] as the backbone, trained and evaluated on the UCF-QNRF dataset. The results are shown in Table I. The SBC framework (i.e., the scenario where no enhancement is used) achieves a mean absolute error (MAE) of 140.6. By utilizing integer-valued bins (E1), the MAE is remarkably reduced from 140.6 to 88.3–a significant 37.1% improvement. This outcome underscores the effectiveness of our approach in reducing ambiguity near the borders. By additionally adopting noise reduction (E2), the MAE further decreases to 85.8. Performance improves even more when the combined loss (E3) is also used. Specifically, increasing the value of $\lambda$ from 0.00 to 1.00 results in improved performance, with the optimal result (77.9) achieved with $\lambda=1.00$.
We also verify that the performance improvement of EBC is agnostic to the choice of backbones. We re-implement three other backbones (ResNet101 [20], MobileNetV2 [21] and DenseNet201 [22]) used by the SBC paper [8], as well as two classic regression-based methods (CSRNet [5] and DMCount [6]) within our EBC framework. The results in Table II show that our EBC framework consistently yields improved performance (up to 44.5%) across different backbones.
IV-B CLIP-EBC Compared with State-of-the-Art
Methods MAE RMSE
P2PNet [13] 72.6 331.6
CLTR [14] 74.4 333.8
STEERER [15] 63.7 309.8
CrowdHat [23] 68.7 296.9
PET [24] 74.4 328.5
CLIP-EBC (ViT-B/16, ours) 61.3 278.4
CLIP-EBC (ViT-L/14, ours) 58.2 268.5
TABLE IV: Comparison of our model CLIP-EBC with state-of-the-art methods on NWPU-Crowd (Test) [11]. CLIP-EBC achieves both the lowest MAE and RMSE on this large-scale benchmark database, surpassing all strong baselines.
We compare CLIP-EBC with the latest crowd-counting methods. Table III presents the results on the ShanghaiTech A & B, UCF-QNRF, and NWPU-Crowd (Val) datasets, while Table  IV illustrates the results on the test split of the NWPU-Crowd database. Our CLIP-EBC models achieve comparable results with state-of-the-art approaches on the ShanghaiTech A & B and UCF-QNRF datasets. On the much larger benchmark database NWPU-Crowd, our CLIP-EBC model with the ViT-B/16 backbone achieves significant performance gains, surpassing all existing models with breakthrough results: an MAE of 36.6 and an RMSE of 81.7 on the validation split. On the test set, the ViT-B/16 backbone achieves an MAE of 61.3 and an RMSE of 278.4, outperforming all strong baselines. Moreover, the ViT-L/14 variant further reduces the errors, achieving an MAE of 58.2 and an RMSE of 268.5, setting new state-of-the-art performance. Fig. 4 provides visualized results on the validation split using CLIP-EBC with the ViT-L/14 backbone.
Figure 4: Visualization of density maps predicted by CLIP-EBC.
V Conclusion
In this paper, we bridged the gap between CLIP and crowd counting by reformulating the counting task as a blockwise classification problem. To this end, we proposed Enhanced Blockwise Classification (EBC), a novel, backbone-agnostic framework that improves upon standard blockwise classification through three key advancements: quantization, noise reduction, and loss formulation. Within our EBC framework, we further proposed CLIP-EBC, the first fully CLIP-based crowd counting method, which successfully leverages CLIP for accurate crowd density estimation. CLIP-EBC classifies local count values into predefined integer-valued bins by comparing the similarity between local image features and text prompt features. Extensive experiments validated the effectiveness of our EBC framework, and the comparison with existing methods on four benchmark datasets demonstrated the superior performance of our CLIP-EBC model.